OpenStudy (anonymous):

3^x=243? how do i figure this out without a calculator?

OpenStudy (zzr0ck3r):

3^2 = 3*3 = 9 3^3 = 3*3*3 = 9*3 = 27 3^4 = 3*3*3*3 = 27*3 = 81 3^5 = 3*3*3*3*3 = 81*3 = 243

OpenStudy (anonymous):

alright :) thanks